# Mafoko: Structuring and Building Open Multilingual Terminologies for South African NLP

### v2

@@ Abstract
Vukosi Marivate1,2,3, Isheanesu Dzingirai1, Fiskani Banda1, Richard Lastrucci1, Thapelo Sindane1, Keabetswe Madumo1, Kayode Olaleye1, Abiodun Modupe1, Unarine Netshifhefhe1, Herkulaas Combrink4,5, Mohlatlego Nakeng1, Matome Ledwaba1 1DSFSI, Dept. of Computer Science, University of Pretoria, 2AfriDSAI, University of Pretoria, 3Lelapa AI, 4Economics and Management Sciences, University of the Free State, 5Interdisciplinary Centre for Digital Futures, University of the Free State Correspondence: vukosi.marivate@cs.up.ac.za
@@ 1 Introduction
The advancement of Natural Language Processing (NLP) is fundamentally tied to the availability of high-quality language resources. However, the vast majority of the world’s languages, including the 12 official languages of South Africa, remain critically under-resourced in this regard (Joshi et al., 2020). This scarcity creates a significant bottleneck for technological development and linguistic preservation. While substantial government and academic initiatives in South Africa have produced multilingual terminology lists over the years (Taljard, 2015), these valuable assets remain largely fragmented, locked in non-machine-readable formats like PDFs, and lack the standardised structure required for modern computational applications.
@@ 2 Motivation
South Africa’s official languages, with the exception of English and to a lesser extent Afrikaans, remain critically under-resourced in the digital domain (Joshi et al., 2020). Despite significant investment from state institutions—including the Department of Sport, Arts and Culture (DSAC), the Pan South African Language Board (PanSALB), and Statistics South Africa (StatsSA), in creating terminologies for crucial domains, these valuable assets are largely unusable for modern NLP. The primary barriers are both technical and legal: resources are frequently published as static, non-machine-readable documents (Handbook, ) and often lack the clear, permissive licensing required for computational reuse and research. This systemic inaccessibility hinders technological development and undermines efforts to achieve linguistic equity in South Africa’s digital sphere.
@@ 3.2 Domain Coverage
The scope of Mafoko is intentionally domain-agnostic, allowing for the inclusion of terminology lists from a wide array of fields. For instance, the DSAC lists encompass domains such as Information and Communication Technology (ICT), Mathematics, Finance, Health Sciences, and Parliamentary Procedure. The StatsSA collection provides comprehensive multilingual terminology for statistics. Similarly, the Open Educational Resource Term Bank (OERTB) project focused on developing African language terminologies for higher education across multiple disciplines (University of Pretoria, 2019; Taljard, 2015). This broad coverage ensures the dataset’s utility across diverse research and application contexts.
@@ 3.3 Challenges in Data Acquisition
A primary challenge was overcoming the fragmented and often inaccessible nature of the source data. This included navigating licensing constraints, which were often unclear or restrictive, and dealing with access limitations, such as portals that only permit single-term queries. The heterogeneity of data formats, ranging from scanned PDFs to structured spreadsheets—required significant and bespoke pre-processing efforts. These hurdles are emblematic of the broader challenges in language resource development for African languages (Taljard et al., 2022). Even within a single source like DSAC, we observed inconsistencies in formatting, such as the representation of parts of speech, across different terminology lists.
@@ 3.4 Data Curation and Structuring
Automated extraction was followed by extensive manual post-processing to ensure the dataset’s quality and utility. A dedicated team member performed detailed cleaning to correct extraction errors, remove artefacts like page headers and garbled characters, and reconstruct table structures to maintain one-to-one alignment between source and target terms. To preserve the authenticity of the original resources, orthographic and formatting variations from the source documents were retained. Where multiple translations existed for a single term, all variants were included to enable the study of lexical variation, synonymy, and regional differences. The statistics of the datasets are available in Table 1
@@
Total Entries 24,226
a English, Afrikaans, and Sepedi (Northern Sotho).
@@ 3.5 Data Release and Availability
The dataset is openly licensed (see Section 3.6) and accessible on multiple platforms, including GitHub, Zenodo, and HuggingFace, to align with FAIR data principles (Wilkinson et al., 2016). We plan to provide both bulk download options and API access. A feedback and validation interface is also under development to enable community-driven refinement by linguists, translators, and other stakeholders. This approach supports a virtuous cycle of continuous improvement, ensuring the resource remains relevant and accurate over time.
@@ 5.1 Task and Models
We evaluated English-to-Tshivenda translation in two distinct domains: Mathematics and Election, using terminology lists from the DSAC. We used two large language models to assess the impact of our RAG approach: the high-performance GPT-4o-mini and the open-source LLaMA3-8B.
@@ 5.2 Experimental Conditions
    RAG with Rare Terms: Terms were selected from the source text based on their low frequency in general English corpora (Reuters, Inaugural Speeches) using the wordfreq library. This strategy focuses the retrieval on the most challenging, domain-specific vocabulary.
@@ 5.3 Evaluation Metrics
We evaluated translation quality using standard automatic metrics: BLEU for n-gram precision, and both chrF and chrF++ for character n-gram recall. Higher scores indicate better translation quality. Results are presented in Table 2.
@@ 5.4.2 Analysis
The results strongly indicate that providing in-context, domain-specific terminology via RAG is a highly effective method for improving LLM translation performance for low-resource languages. The fact that this holds true for both a state-of-the-art proprietary model and a smaller open-source model underscores the robustness of this approach.
@@ 5.5 Discussion
Multilingual Robotics Glossary3 Unisa Accessible (CC BY-NC-SA)
@@
Mechanical Engineering Glossary6 UCT Accessible (PDF)
Economics, Law Glossaries UCT Inaccessible (Named, no links)

### v3

@@ Abstract
Vukosi Marivate1,2,3, Isheanesu Dzingirai1, Fiskani Banda1, Richard Lastrucci1, Thapelo Sindane1, Keabetswe Madumo1, Kayode Olaleye1, Abiodun Modupe1, Unarine Netshifhefhe1, Herkulaas Combrink4,5, Mohlatlego Nakeng1, Matome Ledwaba1 1Data Science for Social Impact, Dept. of Computer Science, University of Pretoria, 2AfriDSAI, University of Pretoria, 3Lelapa AI, 4Economics and Management Sciences, University of the Free State, 5Interdisciplinary Centre for Digital Futures, University of the Free State Correspondence: vukosi.marivate@cs.up.ac.za
@@ 1 Introduction
The advancement of Natural Language Processing (NLP) is fundamentally tied to the availability of high-quality language resources. However, the vast majority of the world’s languages, including the 12 official languages of South Africa, remain critically under-resourced in this regard (Joshi et al., 2020). This scarcity creates a significant bottleneck for technological development and linguistic preservation. While substantial government and academic initiatives in South Africa have produced multilingual terminology lists over the years (Taljard, 2015), these valuable assets remain largely fragmented, locked in non-machine-readable formats like PDFs (sometimes available as scanned PDFs or flattened text without reusable structure), and lack the standardised structure required for modern computational applications.
@@ 2 Motivation
South Africa’s official languages, with the exception of English and to a lesser extent Afrikaans, remain critically under-resourced in the digital domain (Joshi et al., 2020). Despite significant investment from state institutions—including the Department of Sports, Arts and Culture (DSAC), the Pan South African Language Board (PanSALB), and Statistics South Africa (StatsSA), in creating terminologies for crucial domains, these valuable assets are largely unusable for modern NLP. The primary barriers are both technical and legal: resources are frequently published as static, non-machine-readable documents (Handbook, ) and often lack the clear, permissive licensing required for computational reuse and research. This systemic inaccessibility hinders technological development and undermines efforts to achieve linguistic equity in South Africa’s digital sphere. We acknowledge the work of the South African Centre for Digital Language Resources (SADILAR) which has worked to enable creation, preserve some these resources or provide ways to keep track of them (for example through the LwimiLinks [^2] LwimiLinks is a place for multilingual terminology and other useful language resources. https://lwimilinks.sadilar.org project). LwimiLinks still faces the challenge that the data linked or available is mostly in PDF.
@@ 3.2 Domain Coverage
The scope of Mafoko is intentionally domain-agnostic, allowing for the inclusion of terminology lists from a wide array of fields. For instance, the DSAC lists encompass domains such as Information and Communication Technology (ICT), Mathematics, Finance, Health Sciences, and Parliamentary Procedure. The StatsSA collection provides comprehensive multilingual terminology for statistics. Similarly, the Open Educational Resource Term Bank (OERTB) project focused on developing African language terminologies for higher education across multiple disciplines (University of Pretoria, 2019; Taljard, 2015). This broad coverage ensures the dataset’s utility across diverse research and application contexts. We include the University of Pretoria multilingual glossary, the UNISA Multilingual Robotics Glossary: South African Languages Version [^4] https://ir.unisa.ac.za/handle/10500/30440, Multilingual Linguistic Terminology Project (SAMLT) [^5] https://hdl.handle.net/20.500.12185/669 Griesel and Mojapelo (2022) and the AI Terminologies in African Languages in African Languages Contributors (2025).
@@ 3.3 Challenges in Data Acquisition
A primary challenge was overcoming the fragmented and often inaccessible nature of the source data. This included navigating licensing constraints, which were often unclear or restrictive, and dealing with access limitations, such as portals that only permit single-term queries. The heterogeneity of data formats, ranging from scanned PDFs to structured spreadsheets—required significant and bespoke pre-processing efforts. These hurdles are emblematic of the broader challenges in language resource development for African languages (Taljard et al., 2022). Even within a single source like DSAC, we observed inconsistencies in formatting, such as the representation of part-of-speech, across different terminology lists.
@@ 3.4 Data Curation and Structuring
Automated extraction was followed by extensive manual post-processing to ensure the dataset’s quality and utility. A dedicated team member performed detailed cleaning to correct extraction errors, remove artefacts like page headers and garbled characters, and reconstruct table structures to maintain one-to-one alignment between source and target terms. To preserve the authenticity of the original resources, orthographic (use of hyphens, inconsistent capitalisation, or accent marks in lexical entries) and formatting variations from the source documents were retained. Where multiple translations existed for a single term, all variants were included to enable the study of lexical variation, synonymy, and regional differences. The statistics of the datasets are available in Table 1.
@@
Unisa Robotics* Robotics Terminology 11 100
Unisa Multilingual Provide definitions for linguistic terms in nine languages 9 778
UP Glossary Academic Terminology 3 1,768
Total Entries 25,189
* Unisa Robotics glossary is re-released under CC-BY-NC-SA (not compatable with NOODL) but formated like all other Mafoko releases.
@@ 3.5 Data Release and Availability
The dataset is openly licensed (see Section 3.6) and accessible on multiple platforms, including GitHub [^6] https://github.com/dsfsi/za-mafoko, Zenodo [^7] https://doi.org/10.5281/zenodo.17484991 Marivate et al. (2025), and HuggingFace collection [^8] https://huggingface.co/collections/dsfsi/za-mafoko, to align with FAIR data principles (Wilkinson et al., 2016). We are working with SADILAR to contribute to their LwimiLinks initiative (Mafoko is already listed on there), but we also hope to have a mirror of Mafoko available via SADILAR’s digital mirrors. We plan to provide both bulk download options and API access. A feedback and validation interface is also under development to enable community-driven refinement by linguists, translators, and other stakeholders. This approach supports a virtuous cycle of continuous improvement, ensuring the resource remains relevant and accurate over time.
@@ 5.1 Task and Models
We evaluated English-to-Tshivenda translation in two distinct domains: Mathematics and Election, using terminology lists from the DSAC. The datasets that were used for building the RAG terminology were English–Tshivenda Election and Mathematics (Grade R to 6) domain datasets. The election terminology dataset consisted of 576 English–Tshivenda term pairs and their associated context information that defines how a term is used. The mathematics terminology dataset consisted of 966 English–Tshivenda term pairs and their associated context information. The data extraction from PDF files, cleaning and aligning process was automated through a Python script. We used two large language models to assess the impact of our RAG approach: the high-performance GPT-4o-mini and the open-source LLaMA3-8B.
@@ 5.2 Experimental Conditions
    RAG with Rare Terms: Terms were selected from the source text based on their low frequency in general English corpora (Reuters Lewis (1997), Inaugural SpeechesAhrens (2021)) using the wordfreq library. This strategy focuses the retrieval on the most challenging, domain-specific vocabulary.
@@ 5.3 Evaluation Metrics
We evaluated translation quality using 20 elections and 20 mathematics test set pairs using standard automatic metrics: BLEU for n-gram precision, and both chrF and chrF++ for character n-gram recall. Higher scores indicate better translation quality. The test sets were generated using: the 2025/2026 annual teaching plans for grade 4 mathematics for mathematics; the 2016 local government elections post proclamation leaflet and former President Kgalema Motlanthe’S 2009 State of the Nation Address for elections. Results are presented in Table 3 and example sentences used for evaluation are shown in Table 2.
Domain Example Sentences (English and Tshivenda)
Election English: Local government is in your hands!
Tshivenda: Muvhuso wapo u zwanani zwavho!
Mathematics English: The difference between numbers.
Tshivenda: Phambano i re vhukati ha nomboro.
Table 2: Example English and Tshivenda Sentences by Domain
@@ 5.4.2 Analysis
Although the results are low due to the use of early baseline models, the lack of fine-tuning, and a small test set size of 20 samples per domain, the results nonetheless strongly indicate that providing in-context, domain-specific terminology via RAG is a highly effective method for improving LLM translation performance for low-resource languages with the results almost doubling. The fact that this holds true for both a state-of-the-art proprietary model and a smaller open-source model underscores the robustness of this approach.
@@ 5.5 Discussion
LwimiLinks 3 SADILAR Accessible (links to other org resources + PDFs)
@@
Mechanical Engineering and Statistics and Economics Glossaries 6 UCT Accessible (PDF)
@@
CPUT Multilingual Glossraries 10 CPUT Unreliable Access
@@ 8 Limitations
Acknowledgements
The authors gratefully acknowledge the support of the ABSA Chair of Data Science and the Data Science for Social Impact (DSFSI) Lab at the University of Pretoria. This work was supported by UK International Development and the International Development Research Centre (IDRC), Ottawa, Canada, under the AI4D Africa Program. DSFSI also acknowledges gifts from NVIDIA, Google.org, OpenAI, and Meta.